Calculate the solubility (in g/L) of magnesium hydroxide (Ksp = 1.6 × 10-13) in a pH 7.9 buffer. The answer is not .00199 or 15.72
Answer #1

Given: PH=7.9 As we know: PH+POH = 14 POH= 14-PH = 14-7.9 = 6.1 since; POH = -logro [ok] - 6.1= -logro Tour] [on]: 156) = 7.9